OpenStudy (anonymous):

did this on a quiz and got it wrong, does it matter where you start? Solve: -(3x - 2) = -4(x + 1) - 4

OpenStudy (sleepyjess):

Welcome to OpenStudy! First we need to distribute the - and -4

OpenStudy (xapproachesinfinity):

what do you mean by where you start? be specific

OpenStudy (xapproachesinfinity):

\[-(3x-2)=-4(x+1)-4 ~~~like~sleepy~said~distribute -~and~ 4 ~to ~get\] \[-3x+2=-4x-4-4\]

OpenStudy (xapproachesinfinity):

@jspaulding15

OpenStudy (xapproachesinfinity):

@jspaulding115